# Model-free policy gradient for discrete-time mean-field control

### v1

@@ 1 Introduction
Designing algorithms which optimally control large populations of cooperative agents has recently become an important problem due to the breadth and depth of applications. In the face of the complexity of modelling individual interactions between each pair of agents, mean-field control provides a rigorous model particularly suited for problems with many agents interacting in a symmetrical way: instead of considering all possible interactions between every pair of players, one lets the number of agents go to infinity and tries to solve a limiting problem for which it suffices to study the interaction between a typical agent and the probability distribution of its state. An important and practical research direction for such decision-making problems consists in model-free reinforcement learning (RL) algorithms. RL algorithms are based on two elementary approaches: value-based learning algorithms, such as Q-learning [42], where one tries to learn a map which allows to compare the optimal expected reward for different actions and subsequently learn a policy from this map, and policy-based algorithms, also known as policy gradient methods, such as REINFORCE [44], which directly try and optimize the value function with respect to the parameters of a policy. These two approaches are of great importance and are combined in modern deep reinforcement learning algorithmns. The study of such approaches tailored to mean-field control problems has recently received more attention, with many works developing value-based algorithms. However, there is a notable lack of study of model-free policy-based methods for this setting in the literature.
This paper fills this gap by introducing the first policy gradient algorithm for mean-field control problems in discrete-time. As in the classical REINFORCE algorithm for single agent problems, our algorithm is entirely model-free and can be performed by simply observing trajectories of the system.
@@ 1.1 Outline of Main Results
In this section, we provide an informal description of the problem at hand and present our main contributions. We consider a finite-horizon mean-field control problem in discrete time without common noise. We assume that the state space $\mathcal{X}$ is finite and that the action space $\mathcal{A}$ is compact and Polish. Given an admissible control $\alpha=(\alpha_{t})_{t=0}^{T-1}$, the $\mathcal{X}$-valued state follows the dynamics
@@
$\Pi\coloneqq\{\pi:\mathcal{\{}0,\ldots,T-1\}\times\mathcal{X}\times\mathcal{P}(\mathcal{X})\rightarrow\mathcal{P}(\mathcal{A})\text{ measurable}\}.$
@@
$V^{\pi}(\mu_{0})\coloneqq\mathbb{E}^{\pi}_{\xi\sim\mu_{0}}\left[\sum_{t=0}^{T-1}r(X_{t},\alpha_{t},\mathbb{P}_{X_{t}})+g(X_{T},\mathbb{P}_{X_{T}})\right],$ (1.1)
@@
Our goal is to develop a model-free policy gradient algorithm to solve this mean-field control problem. Consider a set of parametrized policies $\{\pi_{\theta}:\theta\in\Theta\}$ with $\Theta\subseteq{\mathbb{R}}^{D}$ given, e.g., by a neural network with trainable parameters. Then we want to find an estimator for $\nabla_{\theta}V^{\pi_{\theta}}(\mu_{0})$ that can easily be computed in a reinforcement learning setting, e.g., by generating trajectories of $X$ with a simulator. In continuous time, this problem has been studied in [11], where the Itô formula along flows of measures is used to write a policy gradient of the value function that can be estimated in a model-free setting. Our main contributions are as follows:
  • We rigorously derive a policy gradient formula for the value function in Proposition 2.1. This result highlights the computational challenges stemming from the mean-field dependence of the transition kernel and the reward functions. In particular, contrary to the single agent problem, it is not possible to directly obtain a model-free estimator of the gradient using a likelihood ratio trick.
@@
  • We present a model-free algorithm to estimate the gradient of the value function, which we call MF-REINFORCE, based on the formula for the gradient of the perturbed value function in Algorithm 1. We provide a detailed analysis and quantitative bounds on the bias and MSE of the newly introduced estimator in Theorems 3.3 and 3.5 respectively, based on the size of the perturbation and number of trajectories sampled.
@@ 1.2 Related Works
Previous works tackle the challenge of learning optimal policies in problems with many agents, both in the mean-field and finite-player settings. The work of [5] investigates policy gradient methods for linear-quadratic MFC in discrete-time. The recent work of [36] studies a (model-based) policy learning algorithm for MFC by simulating a corresponding $N$-particle system. Actor-critic algorithms for asymptotic mean-field games and control problems have been introduced and analyzed in [2, 8]. Surprisingly enough, continuous-time problems have received more attention [9], due to the explicit policy gradient formula one can obtain via Itô’s formula along flows of measures [11], where policy gradient algorithms have been developed using moment neural networks [31], and for linear-quadratic problems [41, 12]. From the perspective of multi-agent reinforcement learning (MARL), [45] uses mean-field symmetry to design policy gradient methods, [29] proposes a natural policy gradient method for heterogeneous problems based on mean-field control. There are many works studying policy gradient algorithms in the context of games, that is when agents compete without a central planner, both in the finite player [22, 15] and mean-field [38, 13, 19, 27] settings.
@@
Many previous works have studied value-based methods, such as $Q$-learning, for MFC. This line of work stems from establishing the dynamic programming principle for the MFC problem [26, 32, 17], or an equivalent mean-field Markov decision process (MFMDP) [30, 4, 6]. Given that equivalent problem, which typically consists of a Markov decision process on augmented state and action spaces, one can define an augmented or integrated $Q$-function and perform $Q$-learning [3, 17, 6], either by discretizing the augmented spaces or using function approximation. This line of work has proven useful to cooperative multi-agent reinforcement learning [16, 18]. Value-based methods have also been explored in continuous-time mean-field control settings [43]. Our approach is different from these previous works since we do not rely on an equivalent MFMDP, and we do not directly learn the optimal value function.
@@
One of the most popular perturbation methods in reinforcement learning consists in adding entropy regularization to the cost functional [46, 10, 34, 21, 14], which has proven effective in mean-field settings [7, 20, 1]. In our work, we take a different approach by directly perturbing the state probability distribution. More generally, perturbation methods are a key concept in model-free / zeroth-order optimization, and such methods have previously been used in a RL context. Early works such as simultaneous perturbation stochastic approximation (SPSA) [37] propose to optimize a loss function by stochastic perturbation of the parameters, an idea further developed in parameter-exploring policy gradient [35] and in the context of black-box policy search [33]. These approaches typically perturb the parameters of the policy, rather than the underlying probability distributions. Distributional perturbations have also appeared in the analysis of robustness in RL [28, 40]. However, to the best of our knowledge, no existing work investigates perturbations of the state-distribution flow itself, nor uses such perturbations to derive a model-free, policy-gradient estimator tailored to mean-field control.
@@ 1.3 Notations and Paper Structure
For a finite space $E$, the space of probability distributions $\mathcal{P}(E)$ is equipped with the total variation distance denoted by $d_{\mathrm{TV}}(\cdot,\cdot)$. For a vector $x=(x_{1},\ldots,x_{D})\in{\mathbb{R}}^{D}$, we write $\|x\|_{\infty}=\sup_{i=1,\ldots,D}|x_{i}|$. For a continuous bounded function $f:\mathcal{E}\rightarrow{\mathbb{R}}^{D}$, where $\mathcal{E}$ is a topological space, we write $\|f\|_{\infty}=\sup_{e\in\mathcal{E}}\|f(e)\|_{\infty}$. For a matrix $A\in{\mathbb{R}}^{m\times n}$, we denote by $\|A\|$ its operator norm induced by the $1$-norm on ${\mathbb{R}}^{m}$ and ${\mathbb{R}}^{n}$, i.e., $\|A\|=\max_{1\leq j\leq n}\sum_{i=1}^{m}|A_{ij}|$. In Section 2, we state the precise working assumptions and present the perturbation technique, we derive the gradient of the perturbed value function and state a convergence result towards the gradient of the true value function. In Section 3, we present the algorithmic details to implement an estimator of the gradient of the perturbed value function and we give quantitative bounds on the bias and mean-squared error (MSE) of the resulting estimator. Section 4 is dedicated to numerical experiments. The proofs of the results stated in Sections 2 and 3 are given in Section 5.
@@ 2.1 Parametrization of State Distributions by Logits
$\operatorname*{softmax}(l)\coloneqq\left(\frac{e^{l_{1}}}{\sum_{i=1}^{d}e^{l_{i}}},\ldots,\frac{e^{l_{d}}}{\sum_{i=1}^{d}e^{l_{i}}}\right)\in{\mathbb{R}}^{d}.$
@@
Hence we can parametrize any distribution on $\mathcal{X}$ with non-zero mass everywhere by a vector of logits in ${\mathbb{R}}^{d}$. Let $\mathcal{P}(\mathcal{X})^{*}\coloneqq\{\mu\in\mathcal{P}(\mathcal{X}):\mu(x)>0,\>\forall x\in\mathcal{X}\}$.
@@
The state space $\mathcal{X}=\{x^{(1)},\ldots,x^{(d)}\}$ is finite and the action space $\mathcal{A}$ is compact and Polish. Moreover, there exist functions $\tilde{r}:\mathcal{X}\times\mathcal{A}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, $\tilde{g}:\mathcal{X}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, $\tilde{P}:\mathcal{X}\times\mathcal{X}\times\mathcal{A}\times{\mathbb{R}}^{d}$, $p:\Theta\times\mathcal{A}\times\mathcal{X}\times\mathcal{P}(\mathcal{X})$, $\tilde{p}:\Theta\times\mathcal{A}\times\mathcal{X}\times{\mathbb{R}}^{d}$ such that
@@
2.2 Policy Gradient Formula: from Single Agent to Mean-Field
Let $V(\mu_{0},\theta)=V^{\pi_{\theta}}(\mu_{0})$. Under Assumption 1, we have $\pi_{\theta}({\mathrm{d}}a\mid t,x,\mu)=p(\theta,a,t,x,\mu)\nu_{\mathcal{A}}({\mathrm{d}}a)$, $\mu_{t}^{\theta}({\mathrm{d}}x)\coloneqq\mathbb{P}(X_{t}^{\theta}=x)\nu_{\mathcal{X}}({\mathrm{d}}x)$, and $l_{t}^{\theta}=\operatorname*{logit}(\mu_{t}^{\theta})$, where $\nu_{\mathcal{X}}$ is the counting measure on $\mathcal{X}$. Consider the joint probability of the whole trajectory $\left((X_{t}^{\theta},\alpha_{t}^{\theta})_{t=0}^{T-1},X_{T}^{\theta})\right)$ under $\pi_{\theta}$, given by
$P_{\theta}({\mathrm{d}}x_{0},{\mathrm{d}}a_{0},\ldots,{\mathrm{d}}a_{T-1},{\mathrm{d}}x_{T})\coloneqq\mu_{0}({\mathrm{d}}x_{0})\prod_{\tau=0}^{T-1}\pi_{\theta}({\mathrm{d}}a_{\tau}\mid\tau,x_{\tau},\mu_{\tau}^{\theta})P({\mathrm{d}}x_{\tau+1}\mid x_{\tau},a_{\tau},\mu_{\tau}^{\theta}).$
@@ 2.2 Policy Gradient Formula: from Single Agent to Mean-Field
In particular, on top of the standard REINFORCE-like term [44], we obtain two additional terms corresponding to the measure derivative of the rewards, and the measure derivative of the transition kernel and policy (due to the mean-field nature of the dynamics).
@@
$\mathrm{RF}(\theta)$ $\coloneqq\int\left(\sum_{t^{\prime}=0}^{T-1}\nabla_{\theta}\log p(\theta,a_{t^{\prime}},t^{\prime},x_{t^{\prime}},\mu_{t^{\prime}}^{\theta})\right)\left(\sum_{t=0}^{T-1}r(x_{t},a_{t},\mu_{t}^{\theta})+g(x_{T},\mu_{T}^{\theta})\right){\mathrm{d}}P_{\theta},$ (2.2)
$\mathrm{MD}(\theta)$ $\coloneqq\int\left(\sum_{t=0}^{T-1}\nabla_{l}\tilde{r}(x_{t},a_{t},l_{t}^{\theta})\nabla_{\theta}l_{t}^{\theta}+\nabla_{l}\tilde{g}(x_{T},l_{T}^{\theta})\nabla_{\theta}l_{T}^{\theta}\right){\mathrm{d}}P_{\theta},$
$\mathrm{MFD}(\theta)$ $\coloneqq\int\left(\sum_{t^{\prime}=0}^{T-1}\nabla_{l}\log\tilde{p}(\theta,a_{t^{\prime}},t^{\prime},x_{t^{\prime}},l_{t^{\prime}}^{\theta})\nabla_{\theta}l_{t^{\prime}}^{\theta}+\nabla_{l}\log\tilde{P}(x_{t^{\prime}+1}\mid x_{t^{\prime}},a_{t^{\prime}},l_{t^{\prime}}^{\theta})\nabla_{\theta}l_{t^{\prime}}^{\theta}\right)$
@@
where $\mathrm{MD(\theta)},\mathrm{MFD}(\theta)$ stand for Measure Derivative and Mean-Field Derivatve respectively.

### v2

@@ 1 Introduction
Designing algorithms which optimally control large populations of cooperative agents has recently become an important problem due to the breadth and depth of applications. In the face of the complexity of modelling individual interactions between each pair of agents, mean-field control provides a rigorous model particularly suited for problems with many agents interacting in a symmetrical way: instead of considering all possible interactions between every pair of players, one lets the number of agents go to infinity and tries to solve a limiting problem for which it suffices to study the interaction between a typical agent and the probability distribution of its state. An important and practical research direction for such decision-making problems consists in model-free reinforcement learning (RL) algorithms. RL algorithms are based on two elementary approaches: value-based learning algorithms, such as $Q$-learning ([42]), where one tries to learn a map which allows one to compare the optimal expected reward for different actions and subsequently learn a policy from this map, and policy-based algorithms, also known as policy gradient methods, such as REINFORCE ([44]), which directly try to optimize the value function with respect to the parameters of a policy. These two approaches are of great importance and are combined in modern deep reinforcement learning algorithms. The study of such approaches tailored to mean-field control problems has recently received more attention, with many works developing value-based algorithms. However, there is a notable lack of study of model-free policy-based methods for this setting in the literature.
This paper fills this gap by introducing the first policy gradient algorithm for mean-field control problems in discrete time. As in the classical REINFORCE algorithm for single-agent problems, our algorithm is entirely model-free and can be performed by simply observing trajectories of the system.
@@ 1.1 Outline of Main Results
In this section, we provide an informal description of the problem at hand and present our main contributions. We consider a mean-field control problem in discrete time without common noise and time horizon $T\in{\mathbb{N}}$. We assume that the state space $\mathcal{X}$ is finite and that the action space $\mathcal{A}$ is compact and Polish. Given an admissible control $\alpha=(\alpha_{t})_{t=0}^{T-1}$, the $\mathcal{X}$-valued state follows the dynamics
@@
$\Pi\coloneq\{\pi:\{0,\ldots,T-1\}\times\mathcal{X}\times\mathcal{P}(\mathcal{X})\rightarrow\mathcal{P}(\mathcal{A})\text{ measurable}\}.$
@@
$V^{\pi}(\mu_{0})\coloneq\mathbb{E}^{\pi}_{\xi\sim\mu_{0}}\left[\sum_{t=0}^{T-1}r(X_{t},\alpha_{t},\mathbb{P}_{X_{t}})+g(X_{T},\mathbb{P}_{X_{T}})\right],$ (1.1)
@@
Our goal is to develop a model-free policy gradient algorithm to solve this mean-field control problem. Consider a set of parametrized policies $\{\pi_{\theta}:\theta\in\Theta\}$ with $\Theta\subseteq{\mathbb{R}}^{D}$ given, e.g., by a neural network with trainable parameters. Then we want to find an estimator for $\nabla_{\theta}V^{\pi_{\theta}}(\mu_{0})$ that can easily be computed in a reinforcement learning setting, e.g., by generating trajectories of $X$ with a simulator. In continuous time, this problem has been studied by [11], where the Itô formula along flows of measures is used to write a policy gradient of the value function that can be estimated in a model-free setting. Our main contributions are as follows:
  • We rigorously derive a policy gradient formula for the value function in Proposition 2.1. This result highlights the computational challenges stemming from the mean-field dependence of the transition kernel and the reward functions. In particular, contrary to the single-agent problem, it is not possible to directly obtain a model-free estimator of the gradient using a likelihood ratio trick.
@@
  • We present a model-free algorithm to estimate the gradient of the value function, which we call MF-REINFORCE, based on the formula for the gradient of the perturbed value function in Algorithm 1. We provide a detailed analysis and quantitative bounds on the bias and mean squared error of the newly introduced estimator in Theorems 3.3 and 3.5 respectively, based on the size of the perturbation and number of trajectories sampled.
@@ 1.2 Related Works
Previous works tackle the challenge of learning optimal policies in problems with many agents, both in the mean-field and finite-player settings. The work of [5] investigates policy gradient methods for linear-quadratic MFC in discrete time. The recent work of [36] studies a (model-based) policy learning algorithm for MFC by simulating a corresponding $N$-particle system. Actor–critic algorithms for asymptotic mean-field games and control problems have been introduced and analyzed by [2] and [8]. Surprisingly enough, continuous-time problems have received more attention (see, e.g., [9]), due to the explicit policy gradient formula one can obtain via Itô’s formula along flows of measures ([11]), where policy gradient algorithms have been developed using moment neural networks ([31]), and for linear-quadratic problems ([41, 12]). From the perspective of multi-agent reinforcement learning (MARL), [45] use mean-field symmetry to design policy gradient methods, [29] propose a natural policy gradient method for heterogeneous problems based on mean-field control. There are many works studying policy gradient algorithms in the context of games, that is when agents compete without a central planner, both in the finite player (see, e.g., [22, 15]) and mean-field (see, e.g., [38, 13, 19, 27]) settings.
@@
Many previous works have studied value-based methods, such as $Q$-learning, for MFC. This line of work stems from establishing the dynamic programming principle for the MFC problem ([26, 32, 17]), or an equivalent mean-field Markov decision process (MFMDP) ([30, 4, 6]). Given that equivalent problem, which typically consists of a Markov decision process on augmented state and action spaces, one can define an augmented or integrated $Q$-function and perform $Q$-learning ([3, 17, 6]), either by discretizing the augmented spaces or using function approximation. This line of work has proven useful to cooperative multi-agent reinforcement learning ([16, 18]). Value-based methods have also been explored in continuous-time mean-field control settings ([43]). Our approach is different from these previous works since we do not rely on an equivalent MFMDP, and we do not directly learn the optimal value function.
@@
One of the most popular perturbation methods in reinforcement learning consists in adding entropy regularization to the cost functional (see, e.g., [46, 10, 34, 21, 14]), which has proven effective in mean-field settings ([7, 20, 1]). In our work, we take a different approach by directly perturbing the state probability distribution. More generally, perturbation methods are a key concept in model-free / zeroth-order optimization, and such methods have previously been used in a RL context. Early works such as simultaneous perturbation stochastic approximation (SPSA) ([37]) propose to optimize a loss function by stochastic perturbation of the parameters, an idea further developed in parameter-exploring policy gradient ([35]) and in the context of black-box policy search ([33]). These approaches typically perturb the parameters of the policy, rather than the underlying probability distributions. Distributional perturbations have also appeared in the analysis of robustness in RL ([28, 40]). However, to the best of our knowledge, no existing work investigates perturbations of the state-distribution flow itself, nor uses such perturbations to derive a model-free, policy-gradient estimator tailored to mean-field control.
@@ 1.3 Notations and Paper Structure
For a compact Polish space $E$, the space of probability distributions $\mathcal{P}(E)$ is equipped with the total variation distance denoted by $d_{\mathrm{TV}}(\cdot,\cdot)$. For a vector $x=(x_{1},\ldots,x_{D})\in{\mathbb{R}}^{D}$, we write $\|x\|_{\infty}=\sup_{i=1,\ldots,D}|x_{i}|$. For a continuous bounded function $f:\mathcal{E}\rightarrow{\mathbb{R}}^{D}$, where $\mathcal{E}$ is a topological space, we write $\|f\|_{\infty}=\sup_{e\in\mathcal{E}}\|f(e)\|_{\infty}$. For a matrix $A\in{\mathbb{R}}^{m\times n}$, we denote by $\|A\|$ its operator norm induced by the $1$-norm on ${\mathbb{R}}^{m}$ and ${\mathbb{R}}^{n}$, i.e., $\|A\|=\max_{1\leq j\leq n}\sum_{i=1}^{m}|A_{ij}|$. In Section 2, we state the precise working assumptions and present the perturbation technique, we derive the gradient of the perturbed value function and state a convergence result towards the gradient of the true value function. In Section 3, we present the algorithmic details to implement an estimator of the gradient of the perturbed value function and we give quantitative bounds on the bias and mean-squared error (MSE) of the resulting estimator. Section 4 is dedicated to numerical experiments. The proofs of the results stated in Sections 2 and 3 are given in Section 5.
@@ 2.1 Parametrization of State Distributions by Logits
$\operatorname*{softmax}(l)\coloneq\left(\frac{e^{l_{1}}}{\sum_{i=1}^{d}e^{l_{i}}},\ldots,\frac{e^{l_{d}}}{\sum_{i=1}^{d}e^{l_{i}}}\right)\in{\mathbb{R}}^{d}.$
@@
Hence we can parametrize any distribution on $\mathcal{X}$ with non-zero mass everywhere by a vector of logits in ${\mathbb{R}}^{d}$. Let $\mathcal{P}(\mathcal{X})^{*}\coloneq\{\mu\in\mathcal{P}(\mathcal{X}):\mu(x)>0,\>\forall x\in\mathcal{X}\}$.
@@
The state space $\mathcal{X}=\{x^{(1)},\ldots,x^{(d)}\}$ is finite and the action space $\mathcal{A}$ is compact and Polish. Moreover, there exist functions $\tilde{r}:\mathcal{X}\times\mathcal{A}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, $\tilde{g}:\mathcal{X}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, $\tilde{P}:\mathcal{X}\times\mathcal{X}\times\mathcal{A}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$, $p:\Theta\times\mathcal{A}\times\{0,\ldots,T-1\}\times\mathcal{X}\times\mathcal{P}(\mathcal{X})\to{\mathbb{R}}$, $\tilde{p}:\Theta\times\mathcal{A}\times\{0,\ldots,T-1\}\times\mathcal{X}\times{\mathbb{R}}^{d}\to{\mathbb{R}}$ such that
@@
2.2 Policy Gradient Formula: from Single-Agent to Mean-Field
Let $V(\mu_{0},\theta)=V^{\pi_{\theta}}(\mu_{0})$. Under Assumption 1, we have $\pi_{\theta}({\mathrm{d}}a\mid t,x,\mu)=p(\theta,a,t,x,\mu)\nu_{\mathcal{A}}({\mathrm{d}}a)$, $\mu_{t}^{\theta}({\mathrm{d}}x)\coloneq\mathbb{P}(X_{t}^{\theta}=x)\nu_{\mathcal{X}}({\mathrm{d}}x)$, and $l_{t}^{\theta}=\operatorname*{logit}(\mu_{t}^{\theta})$, where $\nu_{\mathcal{X}}$ is the counting measure on $\mathcal{X}$. Consider the joint probability of the whole trajectory $\left((X_{t}^{\theta},\alpha_{t}^{\theta})_{t=0}^{T-1},X_{T}^{\theta})\right)$ under $\pi_{\theta}$, given by
$P_{\theta}({\mathrm{d}}x_{0},{\mathrm{d}}a_{0},\ldots,{\mathrm{d}}a_{T-1},{\mathrm{d}}x_{T})\coloneq\mu_{0}({\mathrm{d}}x_{0})\prod_{\tau=0}^{T-1}\pi_{\theta}({\mathrm{d}}a_{\tau}\mid\tau,x_{\tau},\mu_{\tau}^{\theta})P({\mathrm{d}}x_{\tau+1}\mid x_{\tau},a_{\tau},\mu_{\tau}^{\theta}).$
@@ 2.2 Policy Gradient Formula: from Single-Agent to Mean-Field
In particular, on top of the standard REINFORCE-like term ([44]), we obtain two additional terms corresponding to the measure derivative of the rewards, and the measure derivative of the transition kernel and policy (due to the mean-field nature of the dynamics).
@@
$\mathrm{RF}(\theta)$ $\coloneq\int\left(\sum_{t^{\prime}=0}^{T-1}\nabla_{\theta}\log p(\theta,a_{t^{\prime}},t^{\prime},x_{t^{\prime}},\mu_{t^{\prime}}^{\theta})\right)\left(\sum_{t=0}^{T-1}r(x_{t},a_{t},\mu_{t}^{\theta})+g(x_{T},\mu_{T}^{\theta})\right){\mathrm{d}}P_{\theta},$ (2.2)
$\mathrm{MD}(\theta)$ $\coloneq\int\left(\sum_{t=0}^{T-1}\nabla_{l}\tilde{r}(x_{t},a_{t},l_{t}^{\theta})\nabla_{\theta}l_{t}^{\theta}+\nabla_{l}\tilde{g}(x_{T},l_{T}^{\theta})\nabla_{\theta}l_{T}^{\theta}\right){\mathrm{d}}P_{\theta},$
$\mathrm{MFD}(\theta)$ $\coloneq\int\left(\sum_{t^{\prime}=0}^{T-1}\nabla_{l}\log\tilde{p}(\theta,a_{t^{\prime}},t^{\prime},x_{t^{\prime}},l_{t^{\prime}}^{\theta})\nabla_{\theta}l_{t^{\prime}}^{\theta}+\nabla_{l}\log\tilde{P}(x_{t^{\prime}+1}\mid x_{t^{\prime}},a_{t^{\prime}},l_{t^{\prime}}^{\theta})\nabla_{\theta}l_{t^{\prime}}^{\theta}\right)$
@@
where $\mathrm{MD}(\theta),\mathrm{MFD}(\theta)$ stand for Measure Derivative and Mean-Field Derivative respectively.